# History, Development, and Principles of Large Language Models—An Introductory Survey

### v2

@@ 3. PRINCIPLES OF LARGE LANGUAGE MODELS
This section strives to clarify the fundamental principles of LLMs, intending to provide general practitioners with a more profound understanding of the operational mechanisms employed by these models. Fundamental to LLMs is the utilization of deep neural networks to grasp the statistical patterns and semantic nuances of language, enabling the understanding and generation of language. The GPT series, a notable example among these models, incorporates the transformer architecture [vaswani2017attention] and utilizes autoregression to predict consecutive words. The following discussion will concentrate on the GPT-3 [brown2020language] as an exemplar, providing insights into multiple facets that elucidate the principles of LLMs, as illustrated in Figure 3.
@@
Figure 3: The GPT-3 model architecture. Note that only 2 transformer layers are illustrated for simplicity and illustrative purposes and the actual model consists of 12 transformer layers.
@@
  • Embedding: The encoding process above transforms every input word into a one-hot vector of 50,257 dimensions. However, this vector exhibits high sparsity, primarily consisting of zeros, leading to significant space inefficiency. To address this drawback, the subsequent step in the input embedding layer employs a embedding matrix, denoted as $W_{E}$, to condense these 50,257-dimensional input vectors into shorter numeric vectors of length $n$ ($n$ is set to 12,288 for GPT-3). This method condenses information about word meanings into a more compact space, effectively reducing the vector’s length. Specifically, the transformation is represented as: $X_{WordEmbedding}=I_{E}\times W_{E}$, where $X_{WordEmbedding}\in\mathbb{R}^{2048\times 12288}$, $I_{E}\in\mathbb{R}^{2048\times 50257}$ and $W_{E}\in\mathbb{R}^{50257\times 12288}$.
@@
    where $pos$ is the position of the current word within the sentence (e.g. in the sentence “A survey of Large Language Model”, $pos_{A}$=0, $pos_{Large}$=3), $d_{model}$ is the length of the embedding vector 12288 and $i$ is the dimension. However, why delve into such intricate computations for positional encodings? Let’s explore a more straightforward approach to conceptualize it. Initially, considering a text of length $T$, the most basic approach of the positional encoding is counting. This method utilizes $[0,1,2...T-1]$ as positional encodings for each word in the text. For instance, the positional encodings for the $2$-th word would be $[1,1,...,1]$, aligning in length with the embedding vector. However, this encoding strategy presents a notable drawback: within longer sentences, subsequent words hold considerably larger positional encoding values compared to the positional encodings of the initial words. As a result, this disparity inevitably introduces numerical biases in the merged features when combined with word embeddings. To address this issue, normalization becomes a viable consideration. The most straightforward normalization method involves dividing directly by $T$, resulting in positional encodings such as $[0,\frac{1}{T},\ldots,\frac{T-1}{T}]$ (e.g. the 2nd word’s positional encodings=$[\frac{1}{T},\frac{1}{T},\ldots,\frac{1}{T}]$). While this method restricts all positional encodings to the [0,1] range, it introduces a significant inconsistency: the positional values are contingent on the text’s length, causing the positional distinctions between adjacent words in shorter texts to differ from those in longer texts. As illustrated in Figure 4, the positional encodings discrepancy between adjacent word positions in sentence ① is 0.5, while in sentence ②, neighboring word positions are coded with a difference of 0.125. Consequently, neither of these methods proves to be satisfactory. To address this issue, Google devised an alternative approach involving trigonometric functions [vaswani2017attention]. As a result, a positional encoding matrix denoted as $X_{PositionalEncoding}\in\mathbb{R}^{2048\times 12288}$.
@@
    Step 2: Following this, a matrix is derived using the formula $Softmax\leavevmode\nobreak\ (Q\times K^{T})$. This matrix then multiplied with $V$ to obtain the output $Y$, represented as $Y=Softmax\leavevmode\nobreak\ (Q\times K^{T})\times V$, where $Q\in\mathbb{R}^{2048\times 128}$, $K^{T}\in\mathbb{R}^{128\times 2048}$, $V\in\mathbb{R}^{2048\times 128}$, and $Y\in\mathbb{R}^{2048\times 128}$. An illustrative analogy can be drawn to searching for a song within a music software - here, $Q$ denotes the queried song’s name, $K$ encompasses all song names within the software database, and $V$ includes the corresponding audio data of all songs in the software. Thus, $Y$ signifies the audio data of the sought-after song.
@@
    Masked Self-Attention: Masked self-attention is a modified version of self-attention designed to prevent the model from accessing future words during the prediction of the next word by incorporating a masking technique. For example, with the corpus “abcde”, we can train GPT to predict the next word four times by feeding in “a”, “ab”, “abc” and “abcd” separately. However, this involves segmenting and inputting the sentence four times. The mask technique trick streamlines this process, enabling the model to receive the entire sentence at once and generate an output for each word. The challenge arises from the model peeking at subsequent words, which necessitates the use of masking. When predicting the next word of “a” GPT will only see “a”, and when predicting the next word of “ab”, it will only see “ab”. It’s essential to differentiate between self-attention (utilized in BERT [devlin2018bert]) and masked self-attention (utilized in GPT-3 [brown2020language]). Figure 5 provides a visual illustration.
    Multi-Head Self-Attention: Multi-head self-attention subdivides the self-attention mechanism into $h$ independent heads to compute in parallel (for GPT-3, $h$ is 96). Each head focuses on different aspects and connections within the sequence, concatenating their outcomes afterward. This approach facilitates the model to collectively concentrate on information derived from diverse representation subspaces, thereby enhancing its capacity for expression. Each head possesses its distinctive projection matrix, each with dimensions $d=d_{model}/h=12288/96=128$. The outcome from each head is symbolized as $Y_{s}$, and when these individual head outputs are combined through concatenation ($96Y_{s}=12288=96\times 128$), the resultant amalgamation is denoted as $Y$.
  • Feedforward Neural Net: In addition to masked multi-head self-attention sub-layers, each of the transformer layers in the GPT model architecture contains a fully connected feedforward neural net. This neural network serves as a mechanism designed to nonlinearly transform its inputs, thereby enhancing the model’s overall expressive capabilities.
  • Add & Norm: The add & norm modules are incorporated into each transformer layer on the GPT model architecture, where “add” represents the skip connection [kaiming2016deep] and “norm” represents layer normalization [ba2016layer]. The skip connection helps alleviate the issue of gradient vanishing in deep models, enabling the transformer layer to be stacked to a considerable depth. Additionally, layer normalization serves as a regularization strategy to prevent network overfitting. It calculates the mean and variance on each sample independently, without considering other data, thereby mitigating the impact of varying batch sizes.
@@ 4. APPLICATIONS OF LARGE LANGUAGE MODELS
Drug Discovery: The field of drug discovery is characterized by substantial challenges and elevated costs. The utilization of computational methodologies, such as AI, DL, and quantum mechanical techniques, holds the promise of accelerating the identification of potential drug candidates and predicting their properties [sadybekov2023computational, gorgulla2022emerging]. Among recent advancements in this arena, the integration of LLMs, such as ChatGPT, emerges as a valuable addition, adept at generating and analyzing natural language texts [savage2023drug]. LLMs contribute significantly to the initial phase of drug discovery, particularly in identifying suitable drug targets for diverse diseases [haley2020functional, paananen2020omics]. They offer preliminary insights into the structure and functionality of protein-based drug targets, subject to further validation by alternative methodologies. Furthermore, LLMs play a pivotal role in subsequent drug discovery steps, aiding in the design and screening of small molecules capable of interacting with these drug targets [savage2023drug].
@@
Finance: LLMs are driving significant advancements in the finance industry, with significant financial applications including trading and portfolio management [zhang2020deep], financial risk modeling [mashrur2020machine], financial text mining [gupta2020comprehensive, pagliaro2021investor], and financial advisory and customer services [shahfinaid]. AI integration in financial advisory and customer services is an emerging and rapidly expanding field. According to [misischia2022chatbots], AI-driven chatbots already provide more than 37% of supporting functions in various e-service scenarios. Within the financial field, these chatbots are embraced as cost-effective alternatives to human customer service, a trend underscored in the ”Chatbots in Consumer Finance” report [^7] https://www.consumerfinance.gov/data-research/research-reports/chatbots-in-consumer-finance/chatbots-in-consumer-finance. With the advent of LLMs, more and more tasks that were previously considered difficult to handle have become possible, further expanding the potential applications of AI in the financial industry [li2023large]. Notably, there are already some specialized LLMs like BloombergGPT [wu2023bloomberggpt], a 50-billion-parameter model excelling in financial tasks while maintaining strong performance in general language model benchmarks.
@@
Medical: LLMs are now at the forefront of medical AI with immense potential to improve the efficiency and effectiveness of clinical, educational, and research work [thirunavukarasu2023large]. One example is the performance of LLMs on the MedQA dataset [jin2021disease], which serves as a widely used biomedical question-answering dataset consisting of the United States Medical Licensing Examination (USMLE)-style questions. In this dataset, ChatGPT attains human passing levels and [kung2023performance]. Additionally, in a span of fewer than six months, Med-PaLM2 achieves proficiency levels close to human experts [singhal2023towards]. The strong performance of GPT-4 [achiam2023gpt] and Med-PaLM2 in medical tests suggests that LLMs may be useful teaching tools for students currently attaining a lower level in such tests [thirunavukarasu2023large]. GPT-4’s meta-prompt feature allows users to explicitly describe the desired role for the chatbot to take on during conversation; a useful example is a “Socratic tutor mode” [^8] https://sites.highlands.edu/tutorial-center/tutor-resources/online-tutor-training/module-4/the-socratic-method/, which encourages students to think for themselves by pitching questions at decreasing levels of difficulty until students can work out solutions to the fuller question at hand. This mode can foster critical thinking and problem-solving skills among students [thirunavukarasu2023large]. ChatGPT demonstrates superior performance in tasks not necessitating specialized knowledge or when such expertise is available in user prompts [arora2023promise]. For example, consider discharge summaries as a case in point - they entail repetitive tasks involving information interpretation and compression, often requiring minimal specialized knowledge and user prompts [touvron2023llama].
@@
Legal: LLMs exhibit robust capabilities across various legal tasks, particularly in case prediction and generating legal texts. For instance, ChatGPT achieved a 50.3% accuracy rate on NCBE MBE practice exams, significantly surpassing the 25% random guess level, and achieved passing scores in both Evidence and Torts [bommarito2022gpt]. The integration of ChatGPT holds the potential to reshape the entire legal industry. An Australian law firm experimented with ChatGPT to draft a statement of claim based on the 1992 Mabo case, yielding results akin to those of a first-year lawyer [^9] How ChatGPT and other new AI tools are being used by lawyers, architects and coders - ABC News . [iu2023chatgpt] investigation explored the possibility that ChatGPT could replace litigators by evaluating its drafting and research capabilities. While recommending ChatGPT as a complement to litigators rather than a complete replacement, this suggestion stemmed from ChatGPT’s limitations in meeting the precise accuracy and timeliness demands of legal texts, especially in ensuring continuous access to updated legal research. More recently, advancements such as DeLilegal [^10] https://data.delilegal.com/lawQuestion addressed this issue by integrating with external law databases to enhance retrieval and minimize hallucination. Meanwhile, LLMs in the legal field have emerged in recent months such as ChatLaw [cui2023chatlaw].

### v3

@@ 3. PRINCIPLES AND TAXONOMY OF LARGE LANGUAGE MODELS
Fundamental to LLMs is the utilization of deep neural networks to grasp the statistical patterns and semantic nuances of language, enabling the understanding and generation of language. The GPT series, a notable example among these models, incorporates the transformer architecture [77] and utilizes autoregression to predict consecutive words. The following discussion will concentrate on the GPT-3 [10] as an exemplar, providing insights into multiple facets that elucidate the principles of LLMs, as illustrated in Figure 3.
@@
Figure 3: The GPT-3 model architecture. Note that only two transformer layers are illustrated for simplicity and illustrative purposes, and the actual model consists of 12 transformer layers.
@@
  • Embedding: The encoding process above transforms every input word into a one-hot vector of 50,257 dimensions. However, this vector exhibits high sparsity, primarily consisting of zeros, leading to significant space inefficiency. To address this drawback, the subsequent step in the input embedding layer employs an embedding matrix, denoted as $W_{E}$, to condense these 50,257-dimensional input vectors into shorter numeric vectors of length $n$ ($n$ is set to 12,288 for GPT-3). This method condenses information about word meanings into a more compact space, effectively reducing the vector’s length. Specifically, the transformation is represented as: $X_{WordEmbedding}=I_{E}\times W_{E}$, where $X_{WordEmbedding}\in\mathbb{R}^{2048\times 12288}$, $I_{E}\in\mathbb{R}^{2048\times 50257}$ and $W_{E}\in\mathbb{R}^{50257\times 12288}$.
@@
    where $pos$ is the position of the current word within the sentence (e.g., in the sentence “A survey of Large Language Model”, $pos_{A}$=0, $pos_{Large}$=3), $d_{model}$ is the length of the embedding vector 12288 and $i$ is the dimension. However, why delve into such intricate computations for positional encodings? Let’s explore a more straightforward approach to conceptualize it. Initially, considering a text of length $T$, the most basic approach of positional encoding is counting. This method utilizes $[0,1,2...T-1]$ as positional encodings for each word in the text. For instance, the positional encodings for the $2$-th word would be $[1,1,...,1]$, aligning in length with the embedding vector. However, this encoding strategy presents a notable drawback: within longer sentences, subsequent words hold considerably larger positional encoding values compared to the positional encodings of the initial words. As a result, this disparity inevitably introduces numerical biases in the merged features when combined with word embeddings. To address this issue, normalization becomes a viable consideration. The most straightforward normalization method involves dividing directly by $T$, resulting in positional encodings such as $[0,\frac{1}{T},\ldots,\frac{T-1}{T}]$ (e.g., the 2nd word’s positional encodings = $[\frac{1}{T},\frac{1}{T},\ldots,\frac{1}{T}]$). While this method restricts all positional encodings to the [0,1] range, it introduces a significant inconsistency: the positional values are contingent on the text’s length, causing the positional distinctions between adjacent words in shorter texts to differ from those in longer texts. As illustrated in Figure 4, the positional encodings discrepancy between adjacent word positions in sentence ① is 0.5, while in sentence ②, neighboring word positions are coded with a difference of 0.125. Consequently, neither of these methods proves to be satisfactory. To address this issue, Google devised an alternative approach involving trigonometric functions [77]. As a result, a positional encoding matrix denoted as $X_{PositionalEncoding}\in\mathbb{R}^{2048\times 12288}$.
@@
    Step 2: Following this, a matrix is derived using the formula $Softmax~{}(Q\times K^{T})$. This matrix then multiplied with $V$ to obtain the output $Y$, represented as $Y=Softmax~{}(Q\times K^{T})\times V$, where $Q\in\mathbb{R}^{2048\times 128}$, $K^{T}\in\mathbb{R}^{128\times 2048}$, $V\in\mathbb{R}^{2048\times 128}$, and $Y\in\mathbb{R}^{2048\times 128}$. An illustrative analogy can be drawn to searching for a song within a music software - here, $Q$ denotes the queried song’s name, $K$ encompasses all song names within the software database, and $V$ includes the corresponding audio data of all songs in the software. Thus, $Y$ signifies the audio data of the sought-after song.
@@
    Masked Self-Attention: Masked self-attention is a modified version of self-attention designed to prevent the model from accessing future words during the prediction of the next word by incorporating a masking technique. For example, with the corpus “abcde”, we can train GPT to predict the next word four times by feeding in “a”, “ab”, “abc” and “abcd” separately. However, this involves segmenting and inputting the sentence four times. The mask technique trick streamlines this process, enabling the model to receive the entire sentence at once and generate an output for each word. The challenge arises from the model peeking at subsequent words, which necessitates the use of masking. When predicting the next word of “a” GPT will only see “a”, and when predicting the next word of “ab”, it will only see “ab”. It’s essential to differentiate between self-attention (utilized in BERT [71]) and masked self-attention (utilized in GPT-3 [10]). Figure 5 provides a visual illustration.
    Multi-Head Self-Attention: Multi-head self-attention subdivides the self-attention mechanism into $h$ independent heads to compute in parallel (for GPT-3, $h$ is 96). Each head focuses on different aspects and connections within the sequence, concatenating their outcomes afterward. This approach facilitates the model to collectively concentrate on information derived from diverse representation sub-spaces, thereby enhancing its capacity for expression. Each head possesses its distinctive projection matrix, each with dimensions $d=d_{model}/h=12288/96=128$. The outcome from each head is symbolized as $Y_{s}$, and when these individual head outputs are combined through concatenation ($96Y_{s}=12288=96\times 128$), the resultant amalgamation is denoted as $Y$.
  • Feedforward Neural Net: In addition to masked multi-head self-attention sub-layers, each of the transformer layers in the GPT model architecture contains a fully connected feedforward neural net. This neural network serves as a mechanism designed to non-linearly transform its inputs, thereby enhancing the model’s overall expressive capabilities.
  • Add & Norm: The add & norm modules are incorporated into each transformer layer on the GPT model architecture, where “add” represents the skip connection [82] and “norm” represents layer normalization [83]. The skip connection helps alleviate the issue of gradient vanishing in deep models, enabling the transformer layer to be stacked to a considerable depth. Additionally, layer normalization serves as a regularization strategy to prevent network overfitting. It calculates the mean and variance on each sample independently, without considering other data, thereby mitigating the impact of varying batch sizes.
@@
To help readers choose the most suitable model for their specific applications, we further organize LLMs based on their structure. Specifically, we classify LLMs into three types: encoder-only, encoder-decoder, and decoder-only models. This taxonomy is based on the transformer architecture, the backbone of LLMs, comprising core components like the encoder, decoder, and their variations, which together enable their capabilities [77]. Specifically, encoders are designed to process and interpret input sequences, while decoders generate output sequences based on this interpretation [84]. In some models, both components are integrated into an encoder-decoder framework to handle a diverse set of tasks effectively [33]. Conversely, other models focus on specific applications by using either the encoder alone for understanding tasks or the decoder alone for text generation. The detail information about each type of models is as outlined below:
  • Encoder-only models, such as BERT [71], primarily focus on understanding and interpreting input data rather than generating it. They excel in tasks that require strong contextual understanding, such as sentiment analysis, named entity recognition, and text classification. These models are highly effective for analyzing text but are less suitable for text generation tasks due to their limited generation capabilities and computational intensity when handling large texts.
  • Encoder-decoder LLMs, like T5 [62], combine the strengths of both encoders and decoders, making them versatile and powerful for a range of applications. These models are particularly effective for tasks that require both text understanding and generation, such as machine translation, text summarization, and question answering. The encoder-decoder architecture provides a unified approach to these tasks, allowing for a comprehensive understanding and generation of text. However, the complexity of this dual architecture can result in more challenging training processes and slower inference times compared to other model types.
  • Decoder-only LLMs, such as GPT-3 [10], are optimized for generating text, making them ideal for content creation, conversational AI, and creative writing. These models are known for their excellent text-generation capabilities and are effective in few-shot learning scenarios, where they can generate coherent and contextually relevant responses from minimal input. However, decoder-only models can face limitations in understanding context deeply and are often resource-intensive, requiring significant computational power to generate high-quality text outputs.
With the proposed taxonomy, table 1 provides a comparison of state-of-the-art LLMs across different types, with models in each category organized in ascending order of parameter size. This structure enables a systematic examination of key aspects such as architecture, parameter count, provider, release date, open-source status, tunability status, subsequently fine-tuning applied for the model (Adaptation), pre-training data scale, and evaluation methods in their original papers. Analyzing these elements together provides a thorough understanding of the models’ capabilities and differences. Specifically, the type of each model indicates its architecture, whereas the parameter size reflects both the model’s complexity and the resources required for fine-tuning. Next, the provider describes the model’s development context and goals, and the release date places the model within the timeline of technological advancements. Furthermore, open-source status indicates the model’s accessibility for future research and adaptation, whereas tunability status indicates whether the model has been customized after release. Last but not least, the pre-training data scale shows the model’s exposure to various types of information, whereas the evaluation methods assess its performance against benchmarks. As an example, Codex [85], a decoder-only model with 12 billion parameters created by OpenAI and released in July 2021, was pre-trained on 100 billion tokens and supports in-context learning. While it is customizable and open-source, Codex does not incorporate more advanced techniques such as chain-of-thought methods or reinforcement learning with human feedback, which limits its ability to handle certain complex tasks.
@@ 4. APPLICATIONS OF LARGE LANGUAGE MODELS
Table 1: Statistics of large language models in recent years. IT: instruction tuning. RLHF: reinforcement learning with human feedback. ICL: in-context learning. COT: chain-of-thought
Type Model Size Provider Release Open- sourced Tuna- bility Adaptation Pre-train Data scale Evaluation
IT RLHF ICL CoT
Encoder- only ALBERT [81] 0.012 Google AI 02/2020 - - 16GB - -
DeBERTa [29] 0.1 Microsoft 06/2020 - - 80GB - -
ELECTRA [86] 0.11 Google AI 03/2020 - - - - -
BERT [71] 0.11 Google AI 10/2018 - - 3.3B words - -
RoBERTa [87] 0.125 Meta AI 07/2019 - - 160GB - -
XLM-RoBERTa [88] 0.27 Meta AI 07/2020 - - 2.5TB - -
Encoder- decoder ERNIE 3.0 [89] 10 Baidu 07/2021 x - - 300B tokens -
T5 [62] 11 Google AI 10/2019 - - 1T tokens -
T0 [26] 11 BigScience 10/2021 - - -
Flan-T5 [59] 11 Google AI 10/2022 - -
UL2 [27] 20 Google AI 05/2022 - - 1T tokens
AlexaTM [90] 20 Amazon 08/2022 x - - 1.3T tokens
AlphaCode [91] 41 Google AI 02/2022 x x - - 967B tokens - -
Anthropic [92] 52 Anthropic 12/2021 x - - 400B tokens -
NLLB [93] 54.5 Meta AI 07/2022 - - - -
LLaMa [72] 65 Meta AI 02/2023 - - 1.4T tokens -
FLAN [25] 67 Google AI 09/2021 x - - -
Sparrow [94] 70 Google AI 09/2022 x - -
Chinchilla [95] 70 Google AI 03/2022 x - - 1.4T tokens -
OPT-OML [96] 175 Meta AI 12/2022 x - -
Gopher [97] 280 Google AI 12/2021 x - - 300B tokens -
PaLM [60] 540 Google AI 04/2022 - - 780B tokens
U-PaLM [18] 540 Google AI 10/2022 x x - - -
GLaM [31] 1200 Google AI 12/2021 x - - 280B tokens -
Decoder- only Codex [85] 12 OpenAI 07/2021 x x - - 100B tokens -
Pythia [32] 12 Eleuther AI 04/2023 x - - 300B tokens -
CodeGeeX [98] 13 Tsinghua U. 09/2022 - - 850B tokens -
Skywork [99] 13 Kunlun Inc 10/2023 - - 3.2T tokens -
QWEN [100] 14 Alibaba 09/2023 3T tokens -
CodeGen2 [101] 16 Salesforce AI 05/2023 - - 400B tokens -
GPT-NeoX-20B [102] 20 EleutherAI 04/2022 x - - 825GB -
Gemini 1.5 [103] 20 Google AI 02/2024 x - - - -
LlaMa 2 [72] 70 Meta AI 07/2023 2T tokens -
LlaMa 3 [104] 70 Meta AI 04/2024 15T tokens -
HyperCLOVA [17] 82 Naiver 09/2021 x - - 300B tokens -
Galactia [105] 120 Meta AI 11/2022 - - 106B tokens
GLM [28] 130 Tsinghua U. 10/2022 - - 400B tokens -
LaMDA [106] 137 Google AI 01/2022 x - - 768B tokens -
Jurassic-1 [107] 178 AI21 Labs 08/2021 x - - 300B tokens -
GPT-3 [10] 175 OpenAI 05/2020 x x - - 300B tokens -
WebGPT [108] 175 OpenAI 12/2021 x x - - - -
InstructGPT [109] 175 OpenAI 03/2022 x x - -
BLOOM [110] 176 BigScience 11/2022 - - 366B tokens -
BLOOMZ [20] 176 BigScience 11/2022 x - - -
Llama 3.1 [111] 405 Meta AI 06/2024 15T tokens -
GPT-4 [11] 450 OpenAI 03/2023 x 1.76T
MT-NLG [112] 530 Microsoft 01/2022 x - - 967B tokens -
Gemma [113] 2700 Google AI 02/2024 - - 6T tokens -
@@
Software Engineering: LLMs have demonstrated remarkable capabilities across various software engineering tasks such as defect prediction [35], code review [114], code generation [19], and analysis of software-related questions [115, 116, 117]. These models significantly aid in improving code quality, debugging, and reducing human involvement in many aspects of software development. For instance, Codex [85] is a pioneering model in this domain, utilizing a large generative pre-trained model with up to 12 billion parameters to aid developers by generating code, suggesting optimizations, and identifying and correcting errors. Building on this foundation, several other models have been developed to address various facets of coding and software development. Notable examples include DeepMind’s AlphaCode [91], which specializes in generating code for programming competitions, and Meta’s InCoder [23] and Code Llama [118], as well as Salesforce’s CodeRL [119] and CodeGen [24, 101]. The BigCode project has introduced StarCoder [119], and OpenAI has made significant strides with its GPT [11] and ChatGPT series, enhanced through Reinforcement Learning from Human Feedback (RLHF). These developments expand the scope and potential of automated code generation, setting a new standard in software engineering that offers both opportunities and challenges for the industry.
@@
Drug Discovery: The field of drug discovery is characterized by substantial challenges and elevated costs. The utilization of computational methodologies, such as AI, DL, and quantum mechanical techniques, holds the promise of accelerating the identification of potential drug candidates and predicting their properties [120, 121]. Among recent advancements in this arena, the integration of LLMs, such as ChatGPT, emerges as a valuable addition, adept at generating and analyzing natural language texts [122]. LLMs contribute significantly to the initial phase of drug discovery, particularly in identifying suitable drug targets for diverse diseases [123, 124]. They offer preliminary insights into the structure and functionality of protein-based drug targets, subject to further validation by alternative methodologies. Furthermore, LLMs play a pivotal role in subsequent drug discovery steps, aiding in the design and screening of small molecules capable of interacting with these drug targets [122].
@@
Finance: LLMs are driving significant advancements in the finance industry, with significant financial applications including trading and portfolio management [125], financial risk modeling [126], financial text mining [52, 38], and financial advisory and customer services [127]. AI integration in financial advisory and customer services is an emerging and rapidly expanding field. According to [128], AI-driven chatbots already provide more than 37% of supporting functions in various e-service scenarios. Within the financial field, these chatbots are embraced as cost-effective alternatives to human customer service, a trend underscored in the “Chatbots in Consumer Finance” report [^8] https://www.consumerfinance.gov/data-research/research-reports/chatbots-in-consumer-finance/chatbots-in-consumer-finance. With the advent of LLMs, more and more tasks that were previously considered difficult to handle have become possible, further expanding the potential applications of AI in the financial industry [37]. Notably, there are already some specialized LLMs like BloombergGPT [129], a 50-billion-parameter model excelling in financial tasks while maintaining strong performance in general LM benchmarks.
@@
Medical: LLMs are now at the forefront of medical AI with immense potential to improve the efficiency and effectiveness of clinical, educational, and research work [130, 131]. One example is the performance of LLMs on the MedQA dataset [55], which serves as a widely used biomedical question-answering dataset consisting of the United States Medical Licensing Examination (USMLE)-style questions. In this dataset, ChatGPT attains human passing levels and [53]. Additionally, in a span of fewer than six months, Med-PaLM2 achieves proficiency levels close to human experts [132]. The strong performance of GPT-4 [11] and Med-PaLM2 in medical tests suggests that LLMs may be useful teaching tools for students currently attaining a lower level in such tests [130]. GPT-4’s meta-prompt feature allows users to explicitly describe the desired role for the chatbot to take on during conversation; a useful example is a “Socratic tutor mode” [^9] https://sites.highlands.edu/tutorial-center/tutor-resources/online-tutor-training/module-4/the-socratic-method/, which encourages students to think for themselves by pitching questions at decreasing levels of difficulty until students can work out solutions to the fuller question at hand. This mode can foster critical thinking and problem-solving skills among students [130]. ChatGPT demonstrates superior performance in tasks not necessitating specialized knowledge or when such expertise is available in user prompts [133]. For example, consider discharge summaries as a case in point - they entail repetitive tasks involving information interpretation and compression, often requiring minimal specialized knowledge and user prompts [72].
@@
Legal: LLMs exhibit robust capabilities across various legal tasks, particularly in case prediction and generating legal texts. For instance, ChatGPT achieved a 50.3% accuracy rate on NCBE MBE practice exams, significantly surpassing the 25% random guess level, and achieved passing scores in both Evidence and Torts [134]. The integration of ChatGPT holds the potential to reshape the entire legal industry. An Australian law firm experimented with ChatGPT to draft a statement of claim based on the 1992 Mabo case, yielding results akin to those of a first-year lawyer [^10] How ChatGPT and other new AI tools are being used by lawyers, architects and coders - ABC News . [135] investigation explored the possibility that ChatGPT could replace litigators by evaluating its drafting and research capabilities. While recommending ChatGPT as a complement to litigators rather than a complete replacement, this suggestion stemmed from ChatGPT’s limitations in meeting the precise accuracy and timeliness demands of legal texts, especially in ensuring continuous access to updated legal research. More recently, advancements such as DeLilegal [^11] https://data.delilegal.com/lawQuestion addressed this issue by integrating with external law databases to enhance retrieval and minimize hallucination. Meanwhile, LLMs in the legal field have emerged in recent months such as ChatLaw [79].
Education: With the rapid advancement of AI technology, its applications in education are expanding, fundamentally transforming teaching and learning methods. Researchers have recognized the remarkable capabilities of LLMs, such as ChatGPT, and have explored their substantial potential to impact various educational settings. These models can be used in diverse ways, such as simulating students to help train teachers [136, 137, 138] and acting as virtual instructors to provide personalized instruction to students [139, 16, 140]. LLMs can also function as personal tutors, offering real-time feedback [141] and tailored evaluations and suggestions based on individual learning progress [63, 142]. Furthermore, they can enhance the overall learning experience by improving student engagement and promoting greater autonomy in learning [143, 51]. In addition to supporting individualized learning, LLMs can address broader educational challenges, such as the low teacher-student ratio, by supplementing human instructors and providing additional support to students [144]. The integration of LLMs into education presents a unique opportunity to personalize and scale educational resources, making quality education more accessible to a broader range of learners.